OpenStudy (anonymous):

Transform the graph of f(x)=3^x to sketch g(x)=3^−(x+1) −2. Show table of values and each transformation clearly.

OpenStudy (matt101):

Correct me if I'm wrong but the equations we're dealing with are: \[f(x)=3^x\]\[g(x)=3^{-(x+1)}-2\] Do you have any ideas about what the transformations might be?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

flipped over the y-axis shifted left by 1 shifted down by 2 i think those might be the transformations

OpenStudy (matt101):

All correct!

OpenStudy (anonymous):

what would be the table of values?

OpenStudy (matt101):

Use numbers that work easily into the parent equation, in this case f(x)=3^x. So I'd say x-values from 0 to 5 and their corresponding y-values.

OpenStudy (anonymous):